One hundred eight Americans were surveyed to determine the number of hours they spend watching television each month. It was revealed that they watched an average of 151 hours each month with a standard deviation of 25 hours. Assume that the underlying population distribution is normal.

Calculate the EBM.Use a 99% Confidence Level

Round to two decimal places.
Answer #1

Degrees of freedom df = n-1 = 108 - 1 = 107

t critical value at 0.01 significance level for 107 df = 2.623

EBM = t\alpha/2 * S / sqrt(n)

= 2.623 * 25 / sqrt(108)

= 6.31